# Separating Tongue from Thought: Activation Patching Reveals Language-Agnostic Concept Representations in Transformers

### v2

@@ 5 Conclusion
In this paper, we showed that transformers use language-agnostic latent representations of concepts when processing word-level translation prompts. We achieved this by patching latents between parallel forward passes for translation prompts that differed in both input and output languages, as well as in the specific concepts being translated. Our main finding was that translation performance improves when the transformer is forced to translate a concept representation averaged across multiple languages. This finding speaks for language-agnostic concept representations. As we argued, for language-agnostic concept representations, taking the mean representation of a concept across languages should not impair the LLM’s ability to translate this concept. In contrast, for language-specific ones, taking the mean should result in interference between the different language-specific versions of the concept. Thus, our results are consistent with findings from previous work Wendler et al. (2024) indicating that Llama-2 represents concepts in a concept space independent of the language of the prompt. Furthermore, our work also provides evidence that findings from BERT models Wu et al. (2019); Pires et al. (2019) generalize to a wide range of decoder-only transformers.
@@
In this work, we studied how transformers represent concepts when processing multilingual text. However, we only considered very simple translation prompts and also probed only for the language-specific words describing the concept. Further experiments are needed to make claims about how transformers process multilingual text in more general settings. Furthermore, more fine-grained probing will be required to determine to which degree transformers are able to specialize a concept to a language and whether concepts and languages are entangled in more subtle ways.
@@ Appendix B Random prompt task experiment
Appendix C Other models
@@ C.2 Ruling out hypotheses
In addition, instead of just patching in the mean over different language pairs (Figure 10c), we also patch in the mean over contexts composed of different concept words in Figure 10b. In particular, we take the mean over 5 different few-shot contexts from the same language pair. E.g.: Deutsch: “Dorf" - Italiano: “villaggio" Deutsch: “Buch
@@
Our results in Figure 10 show that the mean over contexts does not increase $P(C^{\ell_{T}}_{S})$, whereas the mean over language pairs does. This is intuitive, since there may be some languages in which the mapping from words to concept features results in the correct concept feature vector. Therefore, averaging over different language pairs can increase the signal about the source concept. However, having additional random contexts stemming from the same language pair does not bring in any information about the source concept.
Note that Figure 9 and Figure 10 are on the next two pages.

### v3

@@ 7 Conclusion
In this paper, we showed that transformers use language-agnostic latent representations of concepts when processing word-level translation prompts. We achieved this by patching latents between parallel forward passes for translation prompts that differed in both input and output languages, as well as in the specific concepts being translated. Our main finding was that translation performance improves when the transformer is forced to translate a concept representation averaged across multiple languages. This finding speaks for language-agnostic concept representations. As we argued, for language-agnostic concept representations, taking the mean representation of a concept across languages should not impair the LLM’s ability to translate this concept. In contrast, for language-specific ones, taking the mean should result in interference between the different language-specific versions of the concept. Thus, our results are consistent with findings from previous work Wendler et al. (2024) indicating that Llama-2 represents concepts in a concept space independent of the language of the prompt. Our work also provides evidence that findings from BERT models Wu et al. (2019); Pires et al. (2019) generalize to a wide range of decoder-only transformers.
@@
In this work, we studied how transformers represent concepts when processing multilingual text. However, we only considered very simple concepts, maybe some more complex concepts would have shown a different behavior. Also, we did not study language-specific concepts like “Waldeinsamkeit”, “The feeling of solitude and connectedness to nature when being alone in the woods.”. It would be interesting to see how those are represented.
Furthermore, more fine-grained probing will be required to determine to which degree transformers are able to specialize a concept to a language and whether concepts and languages are entangled in more subtle ways.
@@ Appendix B Random prompt task experiment
Appendix C Other models and languages
@@ C.2 Ruling out hypotheses
In addition, instead of just patching in the mean over different language pairs (Figure 11c, 12c), we also patch in the mean over contexts composed of different concept words in Figure 11b, 12b. In particular, we take the mean over 5 different few-shot contexts from the same language pair. E.g.: Deutsch: “Dorf" - Italiano: “villaggio" Deutsch: “Buch
@@
Our results in Figure 11 and Figure 12 show that the mean over contexts does not increase $P(C^{\ell_{T}}_{S})$, whereas the mean over language pairs does. This is intuitive, since there may be some languages in which the mapping from words to concept features results in the correct concept feature vector. Therefore, averaging over different language pairs can increase the signal about the source concept. However, having additional random contexts stemming from the same language pair does not bring in any information about the source concept.
Note that Figure 9, Figure 11 and Figure 12 are on the next two pages.
We also provide some extra languages in Figure 10
Figure 10: multiple languages with 200 pairs each and 95% confidence interval.
@@
LLama-2 70B
Refer to caption Refer to caption Refer to caption
@@
C.3 Similarity comparison
We experiment with other languages and models in Figure 13 and get the same trends as with our results in Figure 8. We also provide results for another experiment in which instead of measuring embedding similarities, we measure perplexity on ground truth definitions in Figure 14. We did not include this metric in our analysis as it seemed to have less granularity and is more sensitive to the syntax rather than being focused on the semantic.
(a) Aya 23 8B
(b) Gemma 2 2B
Figure 13: Mean similarity between the definition and the mean embedding of the ground truth definitions, as well as the mean similarity between the definition embedding and the embeddings of the definitions of the other concepts in the dataset. For Aya, the results are presented for three target languages: English (with source languages French and German and input language Spanish), French (with source languages Korean, Japanese, Estonian, Finnish and input language English), and Chinese (with source languages Italian, Finnish, Spanish, Russian, Korean and input languages German, Dutch, Chinese, Spanish, Russian). For Gemma, we show English (with source languages Italian, Finnish, Spanish, Russian, Korean and input languages German, Dutch, Chinese, Spanish, Russian), French (with source languages Spanish, German and input language Italian), and Estonian (with source languages English, French, Chinese, German and input language Hindi). We report means and 95% Gaussian confidence intervals computed over the dataset.
Figure 14: Mean loss on the ground truth definitions. We report 95% confidence intervals.